Please help me solve this equation.4r-4 When r=3
Answer by algebrahouse.com(1659)

4r - 4, when r = 3
= 4(3) - 4 {substituted 3 for r}
= 12 - 4 {multiplied}
= 8 {subtracted}

For more help from me, visit: www.algebrahouse.com